# How Many Code and Test Cases Are Enough? Evaluating Test Cases Generation from a Binary-Matrix Perspective

### v1

@@ Abstract
Code evaluation and reinforcement learning rely critically on test cases. However, collecting golden test cases is hard and expensive, motivating the use of LLMs for automatic test case generation. This, in turn, raises a pivotal challenge: how can we rigorously evaluate the quality of the generated test cases? Existing benchmarks often evaluate the exclusion ratio on large, unstructured collections of wrong codes, leading to high computational costs and severe score inflation. Furthermore, they inadvertently reward generators that detect common, trivial bugs, while failing to penalize their inability to identify rare yet critical faults. In this work, we connect two fundamental questions: (1) What is the minimal set of wrong codes sufficient to represent the entire error space? and (2) What is the minimal set of test cases needed to distinguish them? We introduce a novel framework that formalizes benchmark construction as finding an optimal diagnostic basis in a binary code-test matrix, where rows represent wrong codes and columns represent test case results. The rank of this matrix plays a dual role. It specifies the minimal number of independent error patterns, which determines the size of wrong codes. It also provides a tight upper bound on the number of test cases required for complete fault coverage. Our objective is to identify a basis of size equal to the matrix rank that maximizes internal diversity, which is defined as the average pairwise Jaccard similarity of the codes’ failure signatures (i.e., the matrix rows). To tackle this NP-hard problem, we propose WrongSelect, an efficient approximation algorithm combining pre-filtering and random-restart local search to select maximally diverse wrong codes. Applying this framework to millions of competitive programming submissions, we construct TC-Bench, a compact, diverse, and inflation-resistant benchmark. Extensive experiments show that even the most advanced test case generation methods achieve only  60% exclusion rates on TC-Bench, exposing a significant gap in their diagnostic power and highlighting substantial room for future improvement. Our dataset is available at: https://huggingface.co/datasets/Luoberta/TC-Bench and our code is at: https://github.com/Luowaterbi/TC-Bench.
@@ 1 Introduction
To construct our high-quality benchmark, we collect numerous problems with their GTs and user submissions from prestigious algorithm competitions like USACO, NOI, and ICP. We rigorously filter submissions, retaining only those with complete execution results on GTs. Next, we transform the codes for each problem into a binary matrix and calculate its rank to characterize the error pattern complexity. Then, we employ WrongSelect to efficiently select a maximally diverse set of WCs, constructing a structured diagnostic basis (Figure 1 (c) ). Last, we meticulously review, standardize, and translate all problem descriptions into English to ensure consistency and quality. The resulting benchmark, named TC-Bench, contains 877 problems with a total of 9347 WCs. The final set of WCs constitutes less than 2% of the original submissions. This reduction, combined with the principled number of the necessary test cases, can lead to a near-quadratic decrease in evaluation cost, dramatically improving efficiency. To validate TC-Bench, we reproduce and evaluate 5 common test-case generation methods (Jain et al., 2024; Zeng et al., 2025b; Zhang et al., 2023; He et al., 2025; Gu et al., 2024) on 13 leading LLMs (DeepSeek-AI et al., 2024; Int, ; Hui et al., 2024). Experimental results show that even the state-of-the-art method Claude4-Thinking with LCB achieve only approximately 60% performance. By eliminating redundant error patterns and surfacing critical corner cases, TC-Bench ensures that a method’s ability to handle these challenges is directly reflected in its score. This directly prevents the score inflation that plagues less-curated benchmarks.
@@ 2.3 Benchmark Construction
Through this principled pipeline, we ultimately construct TC-Bench, a high-quality diagnostic benchmark with 877 standardized problems, 9347 core WCs, and 7016 correct submissions. More detail are in Appendix B.1.
@@ 3.2 Results
Table 2 presents the results for various model and method combinations on TC-Bench. TC-Bench Reveals a Significant Performance Ceiling for Current Technologies. Even the best-performing combination, Claude-4 + HT, achieves less than 63%. This result strongly validates that WrongSelect indeed selects a diverse and challenging error basis, revealing a performance gap that would otherwise be masked in unfiltered benchmarks. This suggests that there is substantial room for improvement in handling complex and diverse errors, and TC-Bench serves as a reliable yardstick to measure this progress. A High PassRate does not Equate to a High Hackrate, and Performance Gains Primarily Come From WA. A high PassRate score can be hacked by generating a large number of easy test cases. For instance, on Qwen2.5-32B and Deepseek-V3, CRUX’s PassRate is significantly higher than ALGO’s, yet its Hackrate score is substantially lower. The Impact of Methodology Far Outweighs That of the Base Model. The results consistently show that the choice of method has a much greater impact on final performance than the scale or even the source (open-source vs. closed-source) of the base model. For instance, while Qwen2.5-Coder-32B has fewer parameters than the activated parameters of Deepseek-V3, their HackRate scores with the LCB method differ by only 1%. In contrast, on Qwen2.5-Coder-32B, LCB’s HackRate is nearly 40% higher than CRUX. Furthermore, we observe that top-performing open-source models (e.g., Qwen-Coder-Plus) are competitive with leading closed-source models (e.g., the Claude4 series) across various methods. We hypothesize that this is because test case generation is a specialized task that is underrepresented in existing large-scale code pre-training corpora, thus limiting the performance gains through model scaling or a different training corpus. Further experimental analyses, study on Test-Time Scaling and a summary of common error patterns, are detailed in Appendix A.3.
@@ 4 Discussion
Figure 4: (a) shows a comparison of HackRate between all WCs (before filtering) and TC-Bench. (b) The normalized execution time of the correct code. (c) For each random sampling of 8 correct codes, both PassRate and HackRate remain stable.
Unfiltered Code Sets Lead to Severe Score Inflation.
To validate the existence of score inflation, we randomly selected 100 problems and compared the performance of Claude-4-Think on TC-Bench versus on the full set of WCs. Figure 4 (a) is a strong proof. On the full set, LCB exhibits extremely high performance approaching 100%, whereas on TC-Bench its score is just over 50%. Significant performance gaps are also observed on HT and CRUX. This inflation masks the true difficulty of the problems and the deficiencies of the generation methods. A misleadingly perfect score can be achieved by excluding a vast number of simple, repetitive errors, which conceals the method’s incompetence on the few core, difficult error patterns that are the true measure of capability. In contrast, TC-Bench’s systematic filtering, which is grounded in the intrinsic properties of the problem itself, retains a representative and core set of WCs. It thereby corrects for score inflation and serves as a more stable and fair test.
@@
Unlike WCs, which have failure signatures, correct codes all behave identically on GT, differing only in runtime and memory usage. This makes their selection more subtle. Using only a single correct solution as a validator is insufficient. Certain invalid input may still have an output under a specific code. Our initial exploration shows that as the number of correct codes increases(as shown in 9, more ATs are filtered, leading to higher HackRates. However, not all filtering is beneficial. Many complex but valid ATs are wrongly discarded due to timeouts by slow correct codes. Worse, such low-performance correct codes show inconsistency across environments (different OJ platforms). Performance profiling reveals a highly skewed distribution: most correct codes cluster in the top 20% after applying min–max normalization to runtimes (Figure 4 (b)). These high-performance codes are stable across platforms. Consequently, we adopt a biased random sampling strategy: for each problem, we retain only correct codes within the top 20% normalized runtime and randomly sample 8 from this set. Repeated experiments confirm that this strategy yields highly stable evaluation outcomes ( Figure 4 (c)).
@@
An interesting phenomenon emerged during evaluation: some WCs labeled as WA under GTs produce RE or TLE when executed on ATs. To verify whether this is due to server overload, we conduct a controlled experiment. We sample 350 WCs that exhibited RE/TLE and combined them with about 2.6k random WCs. Running these on a 128-core machine, we gradually reduce concurrency from 128 to 88 tasks. The RE/TLE frequency remains nearly constant regardless of system load. This strongly suggests that advanced methods are indeed capable of producing stricter and more challenging ATs than official GTs, revealing hidden bugs related to performance and robustness.
@@ 5 Conclusion
Reproducibility statement
All code referenced in our paper is available at https://anonymous.4open.science/r/TestcaseBenchmark-715C/ and https://anonymous.4open.science/r/TestcaseBench-2A75/. The data processing workflow is described in detail in Section 2 and Appendix B.1.
@@ A.1 Related Work
Evaluation originally followed traditional software testing, emphasizing coverage and distinguishing between buggy and fixed code. (Xu et al., 2025a; Yu et al., 2025c) SWT-Bench (Mündler et al., 2025) and TestGenEval (Jain et al., 2025) transform from SWE-Bench (Jimenez et al., 2024), providing buggy implementations and their corresponding fixes. Others extend beyond single languages or update to recent codebases. For algorithmic problems, TestEval collected 210 problems but still relied on coverage metrics. More recent works shifted toward end-to-end evaluation with large collections of correct and wrong submissions, measuring how often generated test cases exclude incorrect code. (Ma et al., 2025a; Yang et al., 2025; Wang et al., 2025b) However, these approaches either rely on ad-hoc manual selection or expand code sets without selection or analysis. TC-Bench is the first to study how many codes and test cases are sufficient, and provides a principled, efficient evaluation framework.
@@ A.5 Common Failure
When prompted to generate test cases, many LLMs instead output complete solutions to the problem. This issue is particularly common when both test cases and solutions are requested together. We hypothesize that this stems from the infrequency of test-case generation tasks in training data and weakened instruction-following ability after long-cot training (Fu et al., 2025b). DeepSeek-R1 exhibited this issue most severely. Within CRUX and PSEUDO, 74% and 60% of its outputs, respectively, are direct solution code rather than valid test cases. Among the remaining outputs, many are unusable due to formatting errors, such as embedding executable code inside JSON. Because the extractable test cases are too few, R1 is excluded from the main experiments. This finding highlights that successful test-case generation requires not only strong reasoning ability, but also precise task comprehension and robust formatting control.
@@
Many problems require test cases at large boundary conditions. We observe that numerous methods attempt to construct overly large inputs (e.g., huge graph structures), leading to out-of-memory crashes or timeouts during execution. This reveals a deeper limitation: while LLMs are proficient in generating algorithmic logic, they lack awareness of physical execution constraints such as memory and runtime. A robust test-case generation pipeline must therefore incorporate mechanisms like input partitioning or streaming to adapt to limited system resources.

### v2

@@ Abstract
Evaluating test cases automatically generated by Large Language Models (LLMs) is a critical yet challenging task. Existing benchmarks often evaluate the exclusion ratio on large, unstructured collections of wrong codes, suffering from high computational costs and score inflation. Furthermore, they inadvertently reward generators that detect common, trivial bugs, while failing to penalize their inability to identify rare yet critical faults. In this work, we connect two fundamental questions: (1) What is the minimal set of wrong codes sufficient to represent the entire error space? and (2) What is the minimal set of test cases needed to distinguish them? We introduce a novel framework that formalizes benchmark construction as finding an optimal diagnostic basis in a binary code-test matrix, where rows represent wrong codes and columns represent test case results. The rank of this matrix specifies the minimal number of independent error patterns (wrong codes) and provides a tight upper bound on the number of test cases required for complete fault coverage. Our objective is to identify a basis of size equal to the matrix rank that maximizes internal diversity. To tackle this NP-hard problem, we propose WrongSelect, an efficient approximation algorithm to select maximally diverse wrong codes. Applying this framework to millions of competitive programming submissions, we construct TC-Bench, a compact, diverse, and inflation-resistant benchmark. Extensive experiments show that even the most advanced test case generation methods achieve only  60% exclusion rates on TC-Bench, exposing a significant gap in their diagnostic power and highlighting substantial room for future improvement. Our dataset is available at: https://huggingface.co/datasets/Luoberta/TC-Bench and our code is at: https://github.com/Luowaterbi/TC-Bench.
@@ 1 Introduction
To construct our high-quality benchmark, we collect numerous problems with their GTs and user submissions from prestigious algorithm competitions like USACO, NOI, and ICPC. We rigorously filter submissions, retaining only those with complete execution results on GTs. Next, we transform the codes for each problem into a binary matrix and calculate its rank to characterize the error pattern complexity. Then, we employ WrongSelect to efficiently select a maximally diverse set of WCs, constructing a structured diagnostic basis (Figure 1 (c) ). Last, we meticulously review, standardize, and translate all problem descriptions into English to ensure consistency and quality. The resulting benchmark, named TC-Bench, contains 877 problems with a total of 9347 WCs. The final set of WCs constitutes less than 2% of the original submissions. This reduction, combined with the principled number of the necessary test cases, can lead to a near-quadratic decrease in evaluation cost, dramatically improving efficiency. To validate TC-Bench, we reproduce and evaluate 5 common test-case generation methods (Jain et al., 2024; Zeng et al., 2025b; Zhang et al., 2023; He et al., 2025; Gu et al., 2024) on 13 leading LLMs (DeepSeek-AI, A. Liu, B. Feng, B. Xue, B. Wang, B. Wu, C. Lu, C. Zhao, C. Deng, C. Zhang, C. Ruan, D. Dai, D. Guo, D. Yang, D. Chen, D. Ji, E. Li, F. Lin, F. Dai, F. Luo, G. Hao, G. Chen, G. Li, H. Zhang, H. Bao, H. Xu, H. Wang, H. Zhang, H. Ding, H. Xin, H. Gao, H. Li, H. Qu, J. L. Cai, J. Liang, J. Guo, J. Ni, J. Li, J. Wang, J. Chen, J. Chen, J. Yuan, J. Qiu, J. Li, J. Song, K. Dong, K. Hu, K. Gao, K. Guan, K. Huang, K. Yu, L. Wang, L. Zhang, L. Xu, L. Xia, L. Zhao, L. Wang, L. Zhang, M. Li, M. Wang, M. Zhang, M. Zhang, M. Tang, M. Li, N. Tian, P. Huang, P. Wang, P. Zhang, Q. Wang, Q. Zhu, Q. Chen, Q. Du, R. J. Chen, R. L. Jin, R. Ge, R. Zhang, R. Pan, R. Wang, R. Xu, R. Zhang, R. Chen, S. S. Li, S. Lu, S. Zhou, S. Chen, S. Wu, S. Ye, S. Ye, S. Ma, S. Wang, S. Zhou, S. Yu, S. Zhou, S. Pan, T. Wang, T. Yun, T. Pei, T. Sun, W. L. Xiao, W. Zeng, W. Zhao, W. An, W. Liu, W. Liang, W. Gao, W. Yu, W. Zhang, X. Q. Li, X. Jin, X. Wang, X. Bi, X. Liu, X. Wang, X. Shen, X. Chen, X. Zhang, X. Chen, X. Nie, X. Sun, X. Wang, X. Cheng, X. Liu, X. Xie, X. Liu, X. Yu, X. Song, X. Shan, X. Zhou, X. Yang, X. Li, X. Su, X. Lin, Y. K. Li, Y. Q. Wang, Y. X. Wei, Y. X. Zhu, Y. Zhang, Y. Xu, Y. Xu, Y. Huang, Y. Li, Y. Zhao, Y. Sun, Y. Li, Y. Wang, Y. Yu, Y. Zheng, Y. Zhang, Y. Shi, Y. Xiong, Y. He, Y. Tang, Y. Piao, Y. Wang, Y. Tan, Y. Ma, Y. Liu, Y. Guo, Y. Wu, Y. Ou, Y. Zhu, Y. Wang, Y. Gong, Y. Zou, Y. He, Y. Zha, Y. Xiong, Y. Ma, Y. Yan, Y. Luo, Y. You, Y. Liu, Y. Zhou, Z. F. Wu, Z. Z. Ren, Z. Ren, Z. Sha, Z. Fu, Z. Xu, Z. Huang, Z. Zhang, Z. Xie, Z. Zhang, Z. Hao, Z. Gou, Z. Ma, Z. Yan, Z. Shao, Z. Xu, Z. Wu, Z. Zhang, Z. Li, Z. Gu, Z. Zhu, Z. Liu, Z. Li, Z. Xie, Z. Song, Z. Gao, and Z. Pan (2024); 18; B. Hui, J. Yang, Z. Cui, J. Yang, D. Liu, L. Zhang, T. Liu, J. Zhang, B. Yu, K. Lu, K. Dang, Y. Fan, Y. Zhang, A. Yang, R. Men, F. Huang, B. Zheng, Y. Miao, S. Quan, Y. Feng, X. Ren, X. Ren, J. Zhou, and J. Lin (2024)). Experimental results show that even the state-of-the-art method Claude4-Thinking with LCB achieve only approximately 60% performance. By eliminating redundant error patterns and surfacing critical corner cases, TC-Bench ensures that a method’s ability to handle these challenges is directly reflected in its score. This directly prevents the score inflation that plagues less-curated benchmarks.
@@ 2.3 Benchmark Construction
Through this principled pipeline, we ultimately construct TC-Bench, a high-quality diagnostic benchmark with 877 standardized problems, 9347 core WCs, and 7016 correct submissions. More details regarding the construction process are available in Appendix B.1. Furthermore, we present a case study in Appendix C to empirically validate the practical effectiveness of WrongSelect.
@@ 3.2 Results
Table 1 presents the results for various model and method combinations on TC-Bench.
TC-Bench Reveals a Significant Performance Ceiling for Current Technologies. Even the best-performing combination, Claude-4 + HT, achieves less than 63%. This result strongly validates that WrongSelect indeed selects a diverse and challenging error basis, revealing a performance gap that would otherwise be masked in unfiltered benchmarks. This suggests that there is substantial room for improvement in handling complex and diverse errors, and TC-Bench serves as a reliable yardstick to measure this progress.
A High PassRate does not Equate to a High Hackrate. A high PassRate score can be hacked by generating a large number of easy test cases. For instance, on Qwen2.5-32B and Deepseek-V3, CRUX’s PassRate is significantly higher than ALGO’s, yet its Hackrate score is substantially lower.
The Impact of Methodology Far Outweighs That of the Base Model. The results consistently show that the choice of method has a much greater impact on final performance than the scale or even the source (open-source vs. closed-source) of the base model. For instance, while Qwen2.5-Coder-32B has fewer parameters than the activated parameters of Deepseek-V3, their HackRate scores with the LCB method differ by only 1%. In contrast, on Qwen2.5-Coder-32B, LCB’s HackRate is nearly 40% higher than CRUX. Furthermore, we observe that top-performing open-source models (e.g., Qwen-Coder-Plus) are competitive with leading closed-source models (e.g., the Claude4 series) across various methods. We hypothesize that this is because test case generation is a specialized task that is underrepresented in existing large-scale code pre-training corpora, thus limiting the performance gains through model scaling or a different training corpus. Further experimental analyses, study on Test-Time Scaling, and a summary of common error patterns, are detailed in Appendix A.3.
@@ 4 Discussion
Figure 4: (a) Rank distribution of filtered WCs across methods. (b) Comparison of evaluation results against TC-Bench (Ours). (c) Normalized execution times of correct solutions, primarily distributed below $0.2$. (d) Sensitivity analysis showing stable PassRate and HackRate when filtering with random subsets of $8$ correct solutions.
Unfiltered and Heuristic Code Sets Lead to Biased Evaluation.
To validate the impact of code selection strategies, we conduct a rigorous comparison on a subset of 100 problems on Claude-4-Thinking. We compare TC-Bench against three baselines: TCGBench Ma et al. (2025b) (denoted as All_WC), which uses the full set of wrong codes; TestCase-Eval Cao et al. (2025), which randomly samples 20 codes; and TCG Yang et al. (2025), which selects 5 wrong codes from those passing at least 60% of test cases. Unfiltered sets lead to score inflation. As shown in Figure 4 (a), the full set leads to severe score inflation. For instance, LCB exhibits near-perfect performance ($\approx$100%) on All_WC, whereas its score on TC-Bench drops to just over 50%. This inflation masks the method’s incompetence on core, difficult error patterns. Crucially, TestCase-Eval exhibits scores and trends highly similar to All_WC across all methods. This indicates that naive random sampling, while potentially reducing dataset size, fails to exclude redundant error patterns and thus cannot resolve the issue of score inflation. Heuristic Selection results in under-representation. Conversely, TCG yields significantly lower scores. While this might seem rigorous, our rank analysis reveals it stems from insufficient coverage. In Figure 4 (b), the rank of the error space varies significantly per problem. While most are below 20, some approach 30. TCG’s rigid limit of 5 codes forces a drastic under-representation for high-complexity problems. While this lowers the performance scores of current methods, it makes future methods prone to score inflation: they would only need to cover a maximum subset of five patterns rather than the complete error space to achieve perfect scores. In summary, TC-Bench strikes the optimal balance. By selecting a basis defined by the problem’s intrinsic rank, it avoids both the inflation of coverage-based methods and the under-representation of heuristic constraint methods, serving as a stable and fair test suite.
Rank serves as the Upper Bound for the Necessary Number of Test Cases.
The row rank, which represents the number of independent error patterns, equals the column rank, which represents the number of independent diagnostic dimensions. In an error space defined by rank $R$, there are only $R$ linearly independent diagnostic dimensions. Any additional test case is merely a linear combination of these basis dimensions and does not provide new information for distinguishing existing error patterns. Therefore, $R$ test cases are sufficient to distinguish all error patterns, serving as a compact upper bound. Consider a concrete example matrix with $R=3$:
$t_{1}$ $t_{2}$ $t_{3}$ $t_{4}$
$w_{1}$ 1 0 1 0
$w_{2}$ 0 1 1 0
$w_{3}$ 0 1 1 0
$w_{4}$ 0 0 0 1
Here, columns $t_{1}$ and $t_{2}$ are linearly independent, but $t_{3}$ is a linear combination ($t_{3}=t_{1}+t_{2}$). Any wrong code failing on $t_{1}$ or $t_{2}$ implies a predictable behavior on $t_{3}$. Thus, $t_{3}$ offers no new diagnostic dimension. The set $\{t_{1},t_{2},t_{4}\}$ is sufficient to distinguish all unique error patterns. This framework addresses a critical flaw in previous evaluations where the number of test cases was arbitrary. Consequently, problems with small diagnostic dimensions were often “over-tested,” inflating scores, while complex problems were “under-tested.” Using Rank as the budget ensures fairness: it allows simple problems to reveal performance gaps while ensuring complex problems are tested with sufficient depth.
@@
Unlike WCs, which have failure signatures, correct codes all behave identically on GT, differing only in runtime and memory usage. This makes their selection more subtle. Using only a single correct solution as a validator is insufficient. Certain invalid input may still have an output under a specific code. Our initial exploration shows that as the number of correct codes increases (as shown in Figure 8, more ATs are filtered, leading to higher HackRates. However, not all filtering is beneficial. Many complex but valid ATs are wrongly discarded due to timeouts by slow correct codes. Worse, such low-performance correct codes show inconsistency across environments (different OJ platforms). Performance profiling reveals a highly skewed distribution: most correct codes cluster in the top 20% after applying min–max normalization to runtimes (Figure 4 (c)). These high-performance codes are stable across platforms. Consequently, we adopt a biased random sampling strategy: for each problem, we retain only correct codes within the top 20% normalized runtime and randomly sample 8 from this set. Repeated experiments confirm that this strategy yields highly stable evaluation outcomes ( Figure 4 (d)).
@@
An interesting phenomenon emerged during evaluation: some wrong codes labeled as Wrong Answer under GTs produce Runtime Error or Time Limit Error when executed on ATs. To verify whether this is due to server overload, we conduct a controlled experiment. We sample 350 WCs that exhibited RE/TLE and combined them with about 2.6k random WCs. Running these on a 128-core machine, we gradually reduce concurrency from 128 to 88 tasks. The RE/TLE frequency remains nearly constant regardless of system load. This strongly suggests that advanced methods are indeed capable of producing stricter and more challenging ATs than official GTs, revealing hidden bugs related to performance and robustness.
@@ 5 Conclusion
Acknowledge
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004 and 62476073.
@@ A.1 Related Work
Evaluation originally followed traditional software testing, emphasizing coverage and distinguishing between buggy and fixed code. (Xu et al., 2025a; Yu et al., 2025c) SWT-Bench (mündler2025swtbenchtestingvalidatingrealworld) and TestGenEval (Jain et al., 2025) transform from SWE-Bench (Jimenez et al., 2024), providing buggy implementations and their corresponding fixes. Others extend beyond single languages or update to recent codebases. For algorithmic problems, TestEval collected 210 problems but still relied on coverage metrics. More recent works shifted toward end-to-end evaluation with large collections of correct and wrong submissions, measuring how often generated test cases exclude incorrect code. (Ma et al., 2025a; Yang et al., 2025; Wang et al., 2025b) However, these approaches either rely on ad-hoc manual selection or expand code sets without selection or analysis. TC-Bench is the first to study how many codes and test cases are sufficient, and provides a principled, efficient evaluation framework.
Code-Test Matrix
CodeT Chen et al. (2023) and B4 Chen et al. (2024) share the concept of using execution results on test cases (the Code-Test Matrix) as behavioral signatures. CodeT assumes correct code behaviors are consistent while incorrect results are diverse, utilizing signatures for clustering to select a consensus set. B4 calculates the probability of observing the matrix to select the most likely correct cluster. However, these methods aim for solution selection where the correctness of code and tests is unknown, utilizing the matrix primarily for signature matching or probabilistic modeling. In their context, the algebraic rank and basis are not the primary interpretative tools. In contrast, TC-Bench operates on ground truth with guaranteed correct tests and wrong codes. We view the matrix as a complete Error Space and apply linear algebra operations to calculate the Rank and Basis, representing this error space most efficiently.
@@ A.5 Common Failure
When prompted to generate test cases, many LLMs instead output complete solutions to the problem. This issue is particularly common when both test cases and solutions are requested together. We hypothesize that this stems from the infrequency of test-case generation tasks in training data and weakened instruction-following ability after long-cot training (Fu et al., 2025b). DeepSeek-R1 exhibited this issue most severely. As shown in Figure 6, within CRUX and PSEUDO, 74% and 60% of its outputs, respectively, are direct solution code rather than valid test cases. Among the remaining outputs, many are unusable due to formatting errors, such as embedding executable code inside JSON. Because the extractable test cases are too few, R1 is excluded from the main experiments. This finding highlights that successful test-case generation requires not only strong reasoning ability, but also precise task comprehension and robust formatting control.
Figure 6: The figure shows the errors that occur in the direct generation of Testcases by models like R1, using the CRUX and PRESUDO algorithms. Subfigure (a) shows the normal output, subfigure (b) demonstrates the insertion of generated code or direct responses in the form of Testcase code within a string, and subfigure (c) shows the model not following the Testcase generation instructions and instead directly providing the solution.
@@
Many problems require test cases at large boundary conditions. As shown in Figure 7 a, we observe that numerous methods attempt to construct overly large inputs (e.g., huge graph structures), leading to out-of-memory crashes or timeouts during execution. This reveals a deeper limitation: while LLMs are proficient in generating algorithmic logic, they lack awareness of physical execution constraints such as memory and runtime. A robust test-case generation pipeline must therefore incorporate mechanisms like input partitioning or streaming to adapt to limited system resources.
Figure 7: Subfigure a shows that the memory explosion is caused by the model constructing an excessively large number of functions during case generation. Subfigure b, c presents a failed case where the model fails to construct a connected graph as required. The task specifies that all node 1 instances must be able to reach node n, but the constructed graph does not satisfy this connectivity condition.
Failure to Construct Required Complex Data Structures
Some problems in our benchmark admit only test cases with highly constrained structures. As shown in Figure 7 b, c, in one problem, every valid input must be a specific type of connected graph. However, none of the tested methods successfully produced even a single valid input. As a result, these problems ended up with zero usable test cases. This underscores that generating high-difficulty test cases can be as challenging as solving an algorithmic problem, requiring a deep understanding of both data structures and algorithms.
A.6 Results of Supplementary Experiments